OpenStudy (anonymous):

PLEASE HELP ASAP! MEDAL WILL BE REWARDED! Find the second derivative of the algebraic function. f(t)=(9t+2)^2/3

OpenStudy (anonymous):

f'(t)= 18/3 (9t+2)^-(1/3) answer: f''(t)= -18 (9t+2)^-(4/3)